# Large Language Models to Generate System-Level Test Programs Targeting Non-functional Properties

### 15 Mar 2024 v1

@@ Abstract
System-Level Test (SLT) has been a part of the test flow for integrated circuits for over a decade and still gains importance. However, no systematic approaches exist for test program generation, especially targeting non-functional properties of the Device under Test (DUT). Currently, test engineers manually compose test suites from off-the-shelf software, approximating the end-user environment of the DUT. This is a challenging and tedious task that does not guarantee sufficient control over non-functional properties. This paper proposes Large Language Models (LLMs) to generate test programs. We take a first glance at how pre-trained LLMs perform in test program generation to optimize non-functional properties of the DUT. Therefore, we write a prompt to generate C code snippets that maximize the instructions per cycle of a super-scalar, out-of-order architecture in simulation. Additionally, we apply prompt and hyperparameter optimization to achieve the best possible results without further training.
@@ I Introduction
System-Level Test (SLT) is increasingly used to improve the quality assurance of complex Systems-on-Chip. In SLT, the Device under Test (DUT) is placed into an environment that emulates its end-user environment as closely as possible. SLT is executed by running off-the-shelf software in the DUT’s mission mode [1]. As such, SLT programs are expected to exercise execution paths and transactions that are unlikely to be triggered by structural tests, detecting defects that are not covered otherwise [2]. Currently, test engineers manually compose the test suite based on field returns and personal experience together with system engineers. However, this process is challenging and time-consuming, especially given the fact that structural information about the DUT is not available.
For SLT, we consider the DUT to be a black box. This constraint holds since the DUT might contain locked intellectual property, and thus netlists are not available. Additionally, the complexity of the DUT paired with the length of typical SLT programs causes traditional methods, such as automatic test pattern generation and fault simulation, to be infeasible. As such, we cannot use traditional fault models and metrics such as fault coverage to determine the quality of SLT programs.
@@
This paper proposes the use of Large Language Models (LLMs) to automatically generate SLT programs that target non-functional properties of the DUT. The goal is to demonstrate the feasibility of LLMs as code generators in the context of SLT by generating code snippets that maximize the IPC of a super-scalar, out-of-order processor called BOOM [4]. For this purpose, we examine Code Llama [5]. We developed a prompt that tells Code Llama to generate C code and give it some constraints on the output format and contents. Furthermore, we also use prompt and hyperparameter optimization to achieve the best possible results without fine-tuning the LLM.
@@ II Related Work
In this section, we introduce some related work for functional test generation in the context of hardware testing. Afterward, we also list some related work regarding LLMs as an optimizer and for code synthesis.
@@ II-A Test Generation for System-Level Test
There is some work by [6, 7, ] using formal methods to generate small snippets that maximize the switching activity in specific modules in a processor. Their aim is to generate small code snippets for Software-based Self-test using MaxSAT and Genetic Programming. [8, ] introduce a framework based on an automatic test pattern generator and bounded model checking that is able to generate functional tests to detect small delay defects.
In previous work, we have shown that mutation-based greybox fuzzing can be used as a method to generate functional test patterns [9]. The goal of this work was to maximize the power consumption of the BOOM core running on an FPGA.
@@ II-B Large Language Models as Optimizers and Code Generators
Recently, [10, ] proposed to use LLMs as iterative optimizers, e.g., in derivative-free optimization problems. Instead of following previous works for automatic prompt generation [11], the proposed method constructs the prompt using two core pieces of information: previously generated outputs (including corresponding scores) and the optimization problem description. The proposed prompt generation scheme allows creating new prompts in each optimization step and thus iteratively increases the test accuracy based on preceding prompts. Therefore, prompting and using LLMs in a similar way could be highly beneficial if we aim to optimize the non-functional properties of a DUT, for example, maximize instructions per cycle.
@@ III Proposed Approach
Fig. 1 shows the experimental setup. We use the following input prompt: <s>[INST]<<SYS>>You are a C code generator. Only respond with generated code and no explanation. Do not justify the code. Do not return C++ code. Always embed the generated code in Markdown code tags. If a question does not make any sense, or it is not factually coherent, explain why instead of answering something not correct. If you don’t know the answer to a question, please don’t share false information. <</SYS>> Write a single program that aims for a high number of instructions per cycle. Don’t forget to include all the necessary header files. If you use math functions include math.h. [/INST].” The text in between <<SYS>> and <</SYS>> is a so-called system prompt. It sets the context for Code Llama and thus does not need to be repeated for every consecutive user prompt. The other text is called the user prompt and is the actual instruction for Code Llama. The [INST][/INST] pair indicates the complete prompt, as future prompts can contain previous answers outside those to keep a conversation history.
@@
We use prompt-optimizer [13] for prompt optimization. It can be set up to use multiple algorithms, for example, based on entropy (infill low surprising words in the context) or replacing words with synonyms. Fig. 1 shows all available algorithms. We optimize the user and the system prompt separately to use the LLM most efficiently.
@@ IV Experimental Results
The prompt presented in Sec. III has been developed over time from a simple prompt asking Code Llama to generate a program with high IPC since it produced a high number of failing programs. This prompt showed the best results in our experiments. However, we assume that more prompt engineering is required to further improve the program generation.

### 19 Mar 2024 v2

@@ Abstract
System-Level Test (SLT) has been a part of the test flow for integrated circuits for over a decade and still gains importance. However, no systematic approaches exist for test program generation, especially targeting non-functional properties of the Device under Test (DUT). Currently, test engineers manually compose test suites from off-the-shelf software, approximating the end-user environment of the DUT. This challenging and tedious task does not guarantee sufficient control over non-functional properties. This paper proposes Large Language Models (LLMs) to generate test programs. We take a first glance at how pre-trained LLMs perform in test program generation to optimize non-functional properties of the DUT. Therefore, we write a prompt to generate C code snippets that maximize the instructions per cycle of a super-scalar, out-of-order architecture in simulation. Additionally, we apply prompt and hyperparameter optimization to achieve the best possible results without further training.
@@ I Introduction
System-Level Test (SLT) is increasingly used to improve the quality assurance of complex Systems-on-Chip. In SLT, the Device under Test (DUT) is placed into an environment that emulates its end-user environment as closely as possible. SLT is executed by running off-the-shelf software in the DUT’s mission mode [1]. As such, SLT programs are expected to exercise execution paths and transactions unlikely to be triggered by structural tests, detecting defects that are not covered otherwise [2]. Currently, test engineers manually compose the test suite based on field returns and personal experience with system engineers. However, this process is challenging and time-consuming, especially given that structural information about the DUT is unavailable.
For SLT, we consider the DUT to be a black box. This constraint holds since the DUT might contain locked intellectual property, and thus netlists are not available. Additionally, the complexity of the DUT paired with the length of typical SLT programs makes traditional methods, such as automatic test pattern generation and fault simulation, infeasible. As such, we cannot use traditional fault models and metrics such as fault coverage to determine the quality of SLT programs.
@@
This paper proposes using Large Language Models (LLMs) to automatically generate SLT programs that target non-functional properties of the DUT. The goal is to demonstrate the feasibility of LLMs as code generators in SLT by generating code snippets that maximize the IPC of a super-scalar, out-of-order processor called BOOM [4]. For this purpose, we examine Code Llama [5]. We developed a prompt that tells Code Llama to generate C code and gives it some constraints on the output format and contents. Furthermore, we also use prompt and hyperparameter optimization to achieve the best possible results without fine-tuning the LLM.
@@ II Related Work
In this section, we introduce some related work for functional test generation in the context of hardware testing. Afterward, we list some related work regarding LLMs as an optimizer and for code synthesis.
@@ II-A Test Generation for System-Level Test
There is some work by [6, 7, ] using formal methods to generate small snippets that maximize the switching activity in specific modules in a processor. They aim to generate small code snippets for Software-based Self-test using MaxSAT and Genetic Programming. [8, ] introduce a framework based on an automatic test pattern generator and bounded model checking that is able to generate functional tests to detect small delay defects.
In previous work, we have shown that mutation-based greybox fuzzing can be used to generate functional test patterns [9]. This work aimed to maximize the power consumption of the BOOM core running on an FPGA.
@@ II-B Large Language Models as Optimizers and Code Generators
Recently, [10, ] proposed to use LLMs as iterative optimizers, e.g., in derivative-free optimization problems. Instead of following previous works for automatic prompt generation [11], the proposed method constructs the prompt using two core pieces of information: previously generated outputs (including corresponding scores) and the optimization problem description. The proposed prompt generation scheme allows creating of new prompts in each optimization step and thus iteratively increases the test accuracy based on preceding prompts. Therefore, prompting and using LLMs similarly could be highly beneficial if we aim to optimize the non-functional properties of a DUT, such as maximizing instructions per cycle.
@@ III Proposed Approach
Figure 1 shows the experimental setup. We use the following input prompt: <s>[INST]<<SYS>>You are a C code generator. Only respond with generated code and no explanation. Do not justify the code. Do not return C+incorrectays embed the generated code in Markdown code tags. If a question does not make any sense, r it not factually coherent, explain why instead of answering something not correct. If you don’t know the answer to a question, please don’t share false information. <</SYS>> Write a single program that aims for a high number of instructions per cycle. Don’t forget to include all the necessary header files. If you use math functions include math.h. [/INST].” The text in between <<SYS>> and <</SYS>> is a so-called system prompt. It sets the context for Code Llama and thus does not need to be repeated for every consecutive user prompt. The other text is called the user prompt and is the actual instruction for Code Llama. The [INST][/INST] pair indicates the complete prompt, as future prompts can contain previous answers outside those to keep a conversation history.
@@
We use prompt-optimizer [13] for prompt optimization. It can be set up to use multiple algorithms, for example, based on entropy (infill low surprising words in the context) or replacing words with synonyms. Figure 1 shows all available algorithms. We optimize the user and the system prompt separately to use the LLM most efficiently.
@@ IV Experimental Results
The prompt presented in Section III has been developed over time from a simple prompt asking Code Llama to generate a program with high IPC since it produced a high number of failing programs. This prompt showed the best results in our experiments. However, we assume that more prompt engineering is required to further improve the program generation.